# Degrees of Entanglement in Systems of Three Indistinguishable Bosons: Revisiting the GHZ State

### v1

@@ I Introduction
Quantum systems of the same kind can be prepared in states in which their constituents are indistinguishable by any operational means. In other words, exchanging any two particles does not change the composite system’s physical properties. As a consequence, in certain circumstances, the attribution of properties to the individual components of a multipartite quantum system cannot be done in a consistent way. This has deep implications for the analysis of non-classical correlations for systems of multiple bosons or fermions.
@@
Recent theoretical and experimental advances have demonstrated that this indistinguishability can be exploited as a resource for quantum information processing, computation, and metrology [1, 2]. A pivotal aspect of this resource is the entanglement that arises from the exchange symmetry of identical particles. A number of approaches have been put forth with the aim of extracting useful or accessible entanglement from systems whose correlations originate only on the symmetrization of the wave function [1, 3, 4]. Notable developments include the formulation of a resource theory for fermionic systems [5, 6, 2], experimental evidence of entanglement generation in ultracold gases through Bose-Einstein Condensate (BEC) splitting [7], and theoretical studies on the generation of entanglement between molecular condensates [8]. These developments indicate the broad applicability of indistinguishability as a resource across different areas of quantum information science.
@@
Notwithstanding, the study and treatment of the non-classical correlations originated in systems of identical particles still remains challenging. While there exist several efficient tools for dealing with fermionic systems [9, 10, 11, 12, 13], there are still many open questions for the boson case, with the mode entanglement approach [14] (distinct from particle entanglement) being by far the most extensively studied in these systems [15, 16]. Regarding the particle approach, one of the ways in which the separability of systems of indistinguishable particles can be determined, is by analysing their properties as represented by projection operators. The key observation is that, for the distinguishable case, a pure state can be considered separable if and only if it is possible to define a complete set of properties for each of its constituents. A similar criterion can be used for the indistinguishable case, but keeping in mind that it is no longer possible to determine which particle has which property. For a comprehensive review of the properties-based analysis in systems of indistinguishable particles, we refer the reader to [17, 10, 13].
@@
In this work we use the properties-based approach to analyze systems of indistinguishable bosons. We go beyond previous works by studying the bosonic entanglement of systems of three indistinguishable qubits and qutrits. Our work shows that qudit states allow for a larger variety of degrees of entanglement, showing different features than the case of qubit states. Our analysis allows for a classification of the states of these systems in different classes, according to how much separable (or non-separable) they are. We pay special attention to the case of GHZ-like states, and show that these are quite different when we jump from two to three particles. Our method has the advantage that can be implemented numerically, allowing for the characterization of entanglement for higher dimensional models (both in the number of components, and their dimensionality).
@@ II.1 Distinguishable particles
Specifically, for a pure state $|\psi\rangle$ describing a composite two-particle system, the subsystem $S_{1}$ is non-entangled with subsystem $S_{2}$ if there exists a one-dimensional projection operator $\hat{P}$ (with eigenvalue 1, by definition) such that
@@
If this condition holds, we can assert that the first subsystem possesses the physical properties defined by the operator $\hat{P}$. The outcome of measuring $\hat{P}$ is certain, indicating that the first subsystem was prepared in the unique eigenstate of $\hat{P}$.
@@
  • totally entangled, which means that there exist no separable bipartition. This family includes W-like and GHZ-like entangled states that determine two inequivalent classes of entanglement (they cannot be connected by LOCC operations).
@@ II.2 Two indistinguishable particles
Mathematically, and in accordance to 2, we will assert that one of the particles (but we do not know which one) of the pure state $\rho=\ket{\psi}\bra{\psi}$ | has a complete set of properties described by the one-dimensional projection operator $P$ if
@@
The symmetric projector $\mathcal{E}_{P}$ can be interpreted as follows: either one of the particles has the set of properties associated to $P$ and the other does not, or vice versa. With this definition, the condition $Tr[\rho\mathcal{E}_{P}]=1$ implies that at least one of the particles has the property $P$, but we do not know which one [26].
@@ III Three indistinguishable particles
    Given an arbitrary $|\psi\rangle$ check whether there exists $P=|\phi_{0}\rangle\langle\phi_{0}|$ such that $\langle\psi|\mathcal{E}_{P}|\psi\rangle=1$. Notice that this equation can be solved numerically, using a suitable parametrization of the one dimensional projection $P$. If such $P$ does not exists, the state must be considered as completely entangled. If it does, follow the next steps.
@@
with $C_{ijk}$ symmetric under permutations of the indexes. After explicitly separating the first element of the basis, the state takes the form:
@@ III.1 Three indistinguishable qubits
In what follows we present a series of illustrative examples. We shall start/commence with the state W.
@@
At first sight, it appears that the state is totally entangled, similarly as to the distinguishable particles case, since it is not immediately apparent whether any of the particles possess a complete set of properties. However, the behaviour of this state in the indistinguishable regime is different. Moreover, as we will show, the three particle case (the state of Eqn. (19)) is also different than the two particle case (i.e., $|\psi\rangle=\frac{|00\rangle+|11\rangle}{\sqrt{2}}$). In reference [27], the authors demonstrate the existence of one-particle properties for the latter. Upon writing it in the corresponding symmetrized form, it is proven that the state is, in fact, totally separable. However, in the three-particle case, using a numerical implementation of the steps $1-3$ described in Section III we find that at least one particle in (19) possesses the property $P=\ket{\phi_{0}}\bra{\phi_{0}}$ |, with
@@
We can now calculate the desired expectation values $\langle\psi|\mathcal{E}_{P_{i}}|\psi\rangle$, $i=1,2,3$, with $P=|\phi_{0}\rangle\langle\phi_{0}|$ and proceed with the same analysis as before. A straightforward calculation shows that at least one particle has the property $P$, but it is not possible to find other properties orthogonal or equal to $P$. Since $\langle\phi_{0}|\Theta\rangle\neq 0$, $\langle\phi_{0}|\Phi\rangle\neq 0$ and $\langle\Theta|\Phi\rangle\neq 0$, even if the state (19) can be written as a symmetrization of single-particle states, it remains entangled, in contrast to the two-particle case. It is also not maximally entangled, as opposed to the distinguishable GHZ state. In section III.2 we propose an analogous GHZ state for indistinguishable bosonic particles.
@@ III.2.1 GHZ-like states for systems of three qutrits
The above equation means that randomly choosing any two bosons, gives place to a separable state [27], which strengthens the analogy to the distinguishable GHZ state.
@@ IV Conclusions
Our findings verify that non-classical correlations associated with systems of indistinguishable particles are considerably more involved than those observed in distinguishable systems. In particular, bosons exhibit a more complex scenario than fermions, and a systematic approach is still missing. The classification of states according to their separability properties introduced in this work calls for the development of entanglement measures tailored for three-qudits bosonic states. For instance, one might claim that for systems of indistinguishable bosons, at least three particles with a single-particle Hilbert space of dimension greater than two are required in order to recover states with properties analogous to those of the distinguishable qubits GHZ state. A similar phenomenon occurs for identical fermions. In this case, the smallest system exhibiting entanglement is a system comprising two subsystems with a Hilbert space of dimension four and the fermionic GHZ state is obtained for three fermions with a single-particle Hilbert space dimension of six [9, 29, 4] Whether these assertions holds with higher number of particles is an open question which will be addressed in future works.
@@ Acknowledgements.
P.C. and A.P.M. acknowledge funding from Grants No. PICT 2020-SERIEA-00959 from ANPCyT (Argentina) and No. PIP 11220210100963CO from CONICET (Argentina) and partial support from SeCyT, Universidad Nacional de Córdoba (UNC). F.H.H. acknowledges funding from Grants No. PICT 2019-1272 from ANPCyT (Argentina). A.P.M. is grateful to Andrea Valdés-Hernández for fruitful discussions.

### v2

# Degrees of Entanglement in Systems of Three Indistinguishable Bosons: Revisiting the Greenberger-Horne-Zeilinger State

@@ I Introduction
Quantum systems of the same kind can be prepared in states in which their constituents are indistinguishable by any operational means. In other words, exchanging any two particles does not change the composite system’s physical properties. As a consequence, in certain circumstances, the attribution of properties to the individual components of a multipartite quantum system cannot be done in a consistent way. This has deep implications for the analysis of non-classical correlations for systems of multiple bosons or fermions [1, 2].
@@
Recent theoretical and experimental advances have demonstrated that indistinguishability can be exploited as a resource for quantum information processing, computation, and metrology [3, 4, 5, 6, 7, 8, 9, 10]. A pivotal aspect of this resource is the entanglement that arises from the exchange symmetry of identical particles. A number of approaches have been put forth with the aim of extracting useful or accessible entanglement from systems whose correlations originate only on the symmetrization of the wave function [3, 11, 12]. Notable developments include the formulation of a resource theory for fermionic systems [13, 14, 4], experimental evidence of entanglement generation in ultracold gases through Bose-Einstein Condensate (BEC) splitting [15, 16], and theoretical studies on the generation of entanglement between molecular condensates [17]. These developments indicate the broad applicability of indistinguishability as a resource across different areas of quantum information science.
Notwithstanding, the study and treatment of the non-classical correlations originated in systems of identical particles still remains challenging, driving a variety of approaches to address these issues [1, 2, 18, 19, 20, 21, 22, 23, 24, 26, 25].
@@
In particular, while there exist several efficient tools for dealing with fermionic systems [18, 27, 22, 28, 29, 30, 31], there are still many open questions for the boson case, with the mode entanglement approach [32] (distinct from particle entanglement) being by far the most extensively studied in these systems [33, 34]. Regarding the particle approach, one of the ways in which the separability of systems of indistinguishable particles can be determined is by analysing their properties as represented by projection operators. The key observation is that for the distinguishable case a pure state can be considered separable if and only if it is possible to define a complete set of properties for each of its constituents. A similar criterion can be used for the indistinguishable case, but keeping in mind that it is no longer possible to determine which particle has which property. For a comprehensive review of the properties-based analysis in systems of indistinguishable particles, we refer the reader to [1, 27, 29].
@@
In this work we use the properties-based approach to analyze particle entanglement in systems of indistinguishable bosons. We go beyond previous works by studying the entanglement of systems of three indistinguishable qubits and qutrits. Our work shows that qudit states allow for a larger variety of degrees of entanglement, showing different features than the case of qubit states. Our analysis allows for a classification of the states of these systems in different classes, according to how much separable (or non-separable) they are. We pay special attention to the case of GHZ-like states, and show that these are quite different when we jump from two to three particles. Our method has the advantage that it can be implemented numerically, allowing for the characterization of entanglement for higher dimensional models (in both the number of components, and their dimensionality).
@@ II.1 Distinguishable particles
Formulating the previous notion mathematically, it can be stated that in a pure state $|\psi\rangle$ describing a composite two-distinguishable-particle system, the subsystem $S_{1}$ is non-entangled with subsystem $S_{2}$ if there exists a one-dimensional projection operator $\hat{P}$ (with eigenvalue 1, by definition) such that
@@
If this condition holds, we can assert that the first subsystem possesses the physical properties defined by the operator $\hat{P}$ [2]. The outcome of measuring $\hat{P}$ is certain, indicating that the first subsystem was prepared in the unique eigenstate of $\hat{P}$.
@@
  • totally entangled, which means that there exist no separable bipartition. This family includes W-like and GHZ-like entangled states that determine two inequivalent classes of entanglement (they cannot be connected by local operations and classical communication).
@@ II.2 Two indistinguishable particles
Mathematically, and in accordance to (2), we will assert that one of the particles (but we do not know which one) of the pure indistinguishable-particle state $\rho=\ket{\psi}\bra{\psi}$ | has a complete set of properties described by the one-dimensional projection operator $P$ if
@@
Symmetrized observables, such as those in Eq. (5), become relevant in scenarios involving compound systems of identical bosons (or fermions), where correlations arising from the permutational symmetry of the state cannot be neglected. In such cases, it is impossible to consistently assign labels to the particles, and after measurement, one cannot determine which particle is which. Even in the absence of interactions, the permutational symmetry of the state vector can give rise to correlations that have measurable effects. If non-symmetric projections were used in these situations (as in the case of distinguishable particles discussed in Section II.1)), one might erroneously conclude that entanglement is present, even if the particles had never interacted. The use of symmetrized operators allows us to correctly distinguish between correlations resulting from genuine entanglement and those arising solely from permutational symmetry.
The symmetric projector $\mathcal{E}_{P}$ can be interpreted as follows: either one of the particles has the set of properties associated to $P$ and the other does not, or vice versa. With this definition, the condition $Tr[\rho\mathcal{E}_{P}]=1$ implies that one of the particles has the property $P$, but we do not know which one [43, 2].
@@ III Three indistinguishable particles
    Given an arbitrary $|\psi\rangle$ check whether there exists $P=|\phi_{0}\rangle\langle\phi_{0}|$ such that $\langle\psi|\mathcal{E}_{P}|\psi\rangle=1$. Notice that this equation can be solved numerically, using a suitable parametrization of the one dimensional projection $P$. If such $P$ does not exist, the state must be considered as completely entangled. If it does, follow the next steps.
@@
with $C_{ijk}$ symmetric under permutations of the indices. After explicitly separating the first element of the basis, the state takes the form:
@@ III.1 Three indistinguishable qubits
In what follows we present a series of illustrative examples. We shall start with the state W.
@@
At first sight, it appears that the state is totally entangled, as in the case of distinguishable particles, since it is not immediately apparent whether any of the particles possess a complete set of properties and it has the form of the standard distinguishable GHZ state, $|GHZ\rangle=\frac{|000\rangle+|111\rangle}{\sqrt{2}}$. However, the behaviour of this state in the indistinguishable regime is different. Moreover, as we will show, the three particle case (the state of Eqn. (19)) is also different than the two particle case (i.e., $|\psi\rangle=\frac{|00\rangle+|11\rangle}{\sqrt{2}}$). In reference [2], the authors demonstrate the existence of one-particle properties for the latter. Upon writing it in the corresponding symmetrized form, it is proven that the state is, in fact, totally separable. However, in the three-particle case, using a numerical implementation of the steps $1-3$ described in Section III we find that at least one particle in (19) possesses the property $P=\ket{\phi_{0}}\bra{\phi_{0}}$ |, with
@@
We can now calculate the desired expectation values $\langle\psi|\mathcal{E}_{P_{i}}|\psi\rangle$, $i=1,2,3$, with $P=|\phi_{0}\rangle\langle\phi_{0}|$ and proceed with the same analysis as before. A straightforward calculation shows that at least one particle has the property $P$, but it is not possible to find other properties orthogonal or equal to $P$. Since $\langle\phi_{0}|\Theta\rangle\neq 0$, $\langle\phi_{0}|\Phi\rangle\neq 0$ and $\langle\Theta|\Phi\rangle\neq 0$, even if the state (19) can be written as a symmetrization of single-particle states, it remains entangled, in contrast to the two-particle case. It is also not maximally entangled, as opposed to the distinguishable GHZ state. In section III.2 we propose, for the indistinguishable boson case, a state with characteristics analogous of those of the standard GHZ state.
@@ III.2.1 GHZ-like states for systems of three qutrits
The above equation means that randomly choosing any two bosons gives place to a separable state [2], which strengthens the analogy to the distinguishable GHZ state.
@@ IV Conclusions
Our findings verify that non-classical correlations associated with systems of indistinguishable particles are considerably more involved than those observed in distinguishable systems. In particular, bosons exhibit a more complex scenario than fermions, and a systematic approach is still missing. The classification of states according to their separability properties introduced in this work calls for the development of entanglement measures tailored for three-qudit bosonic states. For instance, one might claim that for systems of indistinguishable bosons, at least three particles with a single-particle Hilbert space of dimension greater than two are required in order to recover states with properties analogous to those of the distinguishable qubits GHZ state. A similar phenomenon occurs for identical fermions. In this case, the smallest system exhibiting entanglement is a system comprising two subsystems with a Hilbert space of dimension four and the fermionic GHZ state is obtained for three fermions with a single-particle Hilbert space dimension of six [18, 45, 12]. Whether these assertions holds with higher number of particles is an open question which will be addressed in future works.
@@ Acknowledgements.
P.C. and A.P.M. acknowledge funding from Grants No. PICT 2020-SERIEA-00959 from ANPCyT (Argentina) and No. PIP 11220210100963CO from CONICET (Argentina) and partial support from SeCyT, Universidad Nacional de Córdoba (UNC). F.H.H. acknowledges funding from Grants No. PICT-2019-2019-01272 from ANPCyT (Argentina). A.P.M. is grateful to Andrea Valdés-Hernández for fruitful discussions.